Falconmaster:

Which of the following is the solution set of the given equation? 14 + 8m = 14 - 3m - 5m Ø {0} {all reals}

Falconmaster:

i am thinking either a or b

Falconmaster:

but knowing me i am wrong

Mehek:

First we are going to combine like terms like we did before 14 + 8m is simplified already because neither are like terms however, for 14 - 3m - 5m, -3m and -5m can be combined and -3m - 5m = -8 so 14 + 8m = 14 - 8m

Mehek:

And no matter what value you put for m, it won't equal on both sides because one has + 8m and the other has -8m

Mehek:

Except for 0 14 + 8m = 14 - 8m 14 + 8(0) = 14 - 8(0) 14 + 0 = 14 - 0 14 = 14 So B

Falconmaster:

okay thanks so much

Mehek:

No problem :} Are you understanding this though?

Falconmaster:

yes i am going to write this down in my math notes so i have an example

Mehek:

Oh ok that's good